OpenStudy (anonymous):

Jill Hartman earns $750 per week plus 3 percent of sales in excess of $6,500. If Jill sells $25,000 in 1st week, her earnings are

OpenStudy (ranga):

She earns 3% commission on sales in excess of $6,500. If she sells $25,000 in first week, the excess is 25000 - 6500 = 18500. She earns 3% commission on sales in excess of $6,500. So her commission is 18500 x 3/100 = $555. Add that to her $750 weekly earning to get the total. Total = 750 + 555 = $1,305. If this answer is not there I am quitting!

OpenStudy (anonymous):

that's it thank u

OpenStudy (ranga):

you are welcome.